# A Survey on Federated Fine-Tuning of Large Language Models

### v2

@@ 4.1.2 LoRA in Federated Fine-tuning
FedIT (Zhang et al., 2024b) directly integrates LoRA into the classic FedAvg (McMahan et al., 2017) for instruction tuning. FedSA-LoRA (Guo et al., 2024b) identifies that the $\mathbf{A}$ matrices primarily encode general knowledge, while the $\mathbf{B}$ matrices capture client-specific features; thus, it only uploads $\mathbf{A}$ to the server, significantly reducing communication overhead. FederatedScope-LLM (Kuang et al., 2023) establishes a comprehensive end-to-end pipeline for federated LLM fine-tuning and proposes offsite-tuning strategies to mitigate both communication and computational costs. FeDeRA (Yan et al., 2024) addresses data heterogeneity by initializing LoRA matrices via singular value decomposition on the pre-trained weights. LoRA-FAIR (Bian et al., 2024) introduces a correction mechanism on the server to handle aggregation bias and initialization drift across clients. FLASC (Kuo et al., 2024) incorporates sparsity into LoRA to further reduce communication overhead.
SA-FedLoRA (Yang et al., 2024c) mitigates client drift through parameter regularization and dynamically allocates communication budgets. SLoRA (Babakniya et al., 2023) proposes a novel data-driven initialization scheme to better handle statistical heterogeneity. RoLoRA (Chen et al., 2024d) adopts an alternating minimization approach to improve robustness under non-IID conditions. FedPipe (Fang et al., 2024b) automatically selects critical parameters for fine-tuning and applies quantization to reduce memory usage. LP-FL (Jiang et al., 2023) applies LoRA directly to enable efficient on-device fine-tuning. Fed-piLot (Zhang et al., 2024l) reduces memory consumption through LoRA assignment strategies and introduces a novel spatial-temporal aggregation (STAgg) rule to address heterogeneity. FedRA (Su et al., 2023) adaptively determines parameter update scopes based on client resource constraints, effectively reducing computational, communication, and memory costs.
@@
$\bullet$ Heterogeneous LoRA allows individual clients to adopt different rank values $r$ for their LoRA modules, based on their specific data characteristics or resource constraints. This heterogeneity can manifest either across clients (inter-model) or within different layers of the same model (intra-model). By enabling each client to select a rank that best fits its capabilities and local data, this approach introduces greater flexibility and resource-awareness into the federated fine-tuning process. Table 3 summarizes representative methods and the specific challenges they address.
HETLoRA (Cho et al., 2024) assigns heterogeneous ranks across devices and incorporates rank self-pruning along with sparsity-weighted aggregation to tackle data heterogeneity. FLoRA (Wang et al., 2024f) proposes a stacking-based aggregation scheme and allows devices to select ranks according to their resource budgets. FlexLoRA (Bai et al., 2024a) enables dynamic adjustment of local LoRA ranks to leverage the heterogeneous device resources, while employing singular value decomposition for weight redistribution. LoRA-A2 (Koo et al., 2024) introduces alternating freezing and adaptive rank selection mechanisms to fully utilize heterogeneous device resources while addressing statistical heterogeneity. Byun & Lee (2024) propose a replication-based padding technique to enable aggregation across clients with varying LoRA ranks. FEDHM (Yao et al., 2021) addresses resource constraints by distributing low-rank models with heterogeneous capacities to clients. RBLA (Tavallaie & Nazemi1, ) improves aggregation robustness by simultaneously maintaining and aligning both low-rank and high-rank feature components.
@@
FDLoRA (Qi et al., 2024a) introduces dual LoRA modules on each client to separately capture global and personalized knowledge. Only the global LoRA module are communicated with the central server for aggregating cross-client knowledge. pFedLoRA (Yi et al., 2023) designs a homogeneous small adapter to facilitate federated clients’ heterogeneous local model training, with a proposed iterative training process for global-local knowledge exchange. FedLoRA (Wu et al., 2024f) maintains shared general knowledge in a global full-rank matrix while encoding client-specific knowledge in a personalized low-rank module. FedDPA (Yang et al., 2024b) utilizes a global adapter and a local adapter to jointly address test-time distribution shifts and client-specific personalization. PerFIT (Zhang et al., 2024e) allows each client to search for a personalized architecture by expanding the trainable parameter space of the global model to address data heterogeneity. FEDMEM (Du et al., 2024) equips the global model with a k-nearest neighbor (KNN) classifier that captures client-specific distributional shifts, achieving personalization and overcoming data heterogeneity. FedAMoLE (Zhang et al., 2024i) features an adaptive mixture of LoRA experts (MoLE) module for aggregating heterogeneous models and a reverse selection-based expert assignment strategy that optimizes model architectures based on data distributions.
@@ 5 Datasets and Benchmarks
Table 9: A summary of representative instruction fine-tuning datasets. The construction methods are categorized into three types: human construct, which refers to datasets manually written or annotated by humans; model construct, which refers to data generated by prompting LLMs; and synthetic, which refers to data produced through rule-based or programmatic generation. The datasets span key domains such as general language understanding, finance, medicine, code, math, and law.
Dataset Language Construction Method Domain Description
Alpaca English Model Construct General Generated by Text-Davinci-003 with Alpaca-style instruction prompts.
Alpaca-GPT4 English Model Construct General Generated by GPT-4 based on Alpaca prompts with more nuanced multi-turn instructions.
Self-Instruct English Human + Model Construct General Dataset from seed instructions using GPT-3 for improving model generalization.
UltraChat 200k English Model Construct General High-quality multi-turn dialogue dataset filtered from UltraChat.
OpenOrca English Model Construct General Dataset of $\sim$ 4.2M GPT-3.5/4 augmented FLAN examples for instruction following.
ShareGPT90K English Model Construct General Multi-turn dialogue dataset of 90K samples derived from ShareGPT for instruction following.
WizardLM Evol-Instruct V2 196k English Model Construct General Dataset of 196K samples generated via Evol-Instruct method.
Databricks Dolly 15K English Human Construct General Dataset of 15K human-generated prompt-response pairs across diverse tasks.
Baize English Model Construct General Instruction-following dialogues generated by prompting ChatGPT with user-centric queries.
OpenChat English Model Construct General Instruction-following dialogues generated using open LLMs for multi-turn alignment.
Flan-v2 English Model Construct General Dataset combining Flan, P3, Super-Natural Instructions, Chain-of-thought, and Dialog tasks.
BELLE-train-0.5M-CN Chinese Human + Model Construct General Dataset of 519K Chinese samples for instruction following.
BELLE-train-1M-CN Chinese Human + Model Construct General Dataset of 917K Chinese samples for instruction following.
BELLE-train-2M-CN Chinese Human + Model Construct General Dataset of 2M Chinese samples for instruction following.
Firefly-train-1.1M Chinese Human Construct General Dataset of 1.65M Chinese samples across 23 tasks with human-written instruction templates.
Wizard-LM-Chinese-instruct-evol Chinese Human + Model Construct General Dataset of 70K samples translated from WizardLM to Chinese with GPT-generated responses.
HC3-Chinese Chinese Human + Model Construct General Chinese Human-ChatGPT QA pairs across multiple domains.
HC3 English / Chinese Human + Model Construct General Human-ChatGPT QA pairs across multiple domains.
ShareGPT-Chinese-English-90k English / Chinese Model Construct General Bilingual human-machine QA dataset of 90K real user queries in Chinese and English.
FinGPT English Human + Model Construct Finance Instruction tuning data for financial tasks.
Finance-Instruct-500k English Human + Model Construct Finance Large-scale instruction tuning dataset for financial tasks.
Finance-Alpaca English Human + Model Construct Finance Instruction tuning dataset combining Alpaca and FiQA.
Financial PhraseBank English Human Construct Finance Manually annotated financial news sentences for sentiment classification.
Yahoo-Finance-Data English Human Construct Finance Financial dataset including estimates, statements, and historical prices from Yahoo Finance.
Financial-QA-10K English Model Construct Finance Contextual QA dataset for financial question answering and retrieval tasks.
Financial-Classification English Human Construct Finance Dataset combining Financial PhraseBank and Kaggle financial texts.
Twitter-Financial-News-Topic English Human Construct Finance Annotated tweets for multi-class financial topic classification tasks.
Financial-News-Articles English Human Construct Finance Financial news articles for text classification and sentiment analysis tasks.
FiQA English Human Construct Finance QA dataset from financial texts and forums covering finance topics.
Earnings-Call English Human Construct Finance QA pairs from CEO/CFO earnings calls for financial reasoning.
Doc2EDAG Chinese Human Construct Finance Financial reports annotated for event graph extraction and structuring.
Synthetic-PII-Finance-Multilingual Multilingual Synthetic Finance Financial documents with labeled PII for NER and privacy-preserving model training.
ChatDoctor-200K English Human + Model Construct Medical Instruction tuning dataset for medical QA and dialogue generation.
ChatDoctor-HealthCareMagic-100k English Human Construct Medical Real-world doctor-patient conversations.
Medical Meadow CORD-19 English Human Construct Medical Instruction tuning dataset for medical literature summarization based on CORD-19.
Medical Meadow MedQA English Human Construct Medical Multiple-choice medical QA dataset derived from professional medical board exams.
HealthCareMagic-100k-en English Human Construct Medical Real doctor-patient consultations for instruction tuning.
ChatMed-Consult-Dataset Chinese Human + Model Construct Medical Instruction tuning dataset of Chinese medical consultations for QA and dialogue generation.
CMtMedQA Chinese Human Construct Medical Multi-turn medical QA dataset with real doctor-patient conversations.
DISC-Med-SFT Chinese Human + Model Construct Medical Instruction tuning dataset combining real-world dialogues and knowledge graph QA pairs.
Huatuo-26M Chinese Human Construct Medical Structured QA pairs extracted from medical encyclopedias and articles.
Huatuo26M-Lite Chinese Human + Model Construct Medical Refined subset of Huatuo-26M with ChatGPT-rewritten answers.
ShenNong-TCM-Dataset Chinese Human + Model Construct Medical QA pairs generated via entity-centric self-instruct from a TCM knowledge graph.
HuatuoGPT-sft-data-v1 Chinese Human + Model Construct Medical Instruction tuning dataset combining ChatGPT-generated and real doctor-patient dialogues.
MedDialog English / Chinese Human Construct Medical Large-scale doctor-patient dialogues for medical QA and dialogue generation.
CodeAlpaca English Model Construct Code GPT-generated code instruction-following dataset based on the Alpaca style.
Code Instructions 120k Alpaca English Human + Model Construct Code Instruction tuning dataset for code generation with prompts in alpaca style.
CodeContests English Human Construct Code Dataset from competitive programming platforms, designed for program synthesis and reasoning.
CommitPackFT English Human Construct Code Filtered dataset of GitHub commits with high-quality messages for code generation tasks.
ToolBench English Human + Model Construct Code Instruction tuning dataset for multi-tool API usage scenarios.
CodeParrot-Clean English Human Construct Code Deduplicated and filtered dataset of Python files from GitHub for code generation.
The Stack v2 Dedup English Human Construct Code Large-scale deduplicated dataset of source code from 600+ programming languages.
CodeSearchNet English Human Construct Code Code-language pairs for retrieval and semantic search across six programming languages.
CodeForces-CoTs English Human + Model Construct Code A dataset consists of 10k CodeForces problems with reasoning traces generated by DeepSeek R1.
CodeXGLUE Code Refinement English Human Construct Code Dataset of buggy and fixed Java functions for code refinement.
GSM8K English Human Construct Math A dataset of 8.5k grade school-level arithmetic word problems with detailed step-by-step solutions.
CoT-GSM8k English Human + Model Construct Math Extended version of GSM8K with chain-of-thought reasoning steps.
MathInstruct English Human + Model Construct Math Dataset combining CoT and program-of-thought rationales across diverse mathematical fields.
MetaMathQA English Human + Model Construct Math Multi-perspective question augmentations from GSM8K and MATH.
OpenR1-Math-220k English Human + Model Construct Math Dataset of 220k math problems with multiple reasoning traces generated by DeepSeek R1.
Hendrycks MATH Benchmark English Human Construct Math Dataset of 12.5K high school competition math problems with step-by-step solutions.
DeepMind Mathematics Dataset English Synthetic Math Dataset of algorithmically generated math problems across various topics.
OpenMathInstruct-1 English Human + Model Construct Math Dataset of 1.8M math problems with code-interpreter solutions generated by Mixtral-8x7B.
Orca-Math Word Problems 200k English Synthetic Math Synthetic dataset of 200K grade school math word problems with GPT-4 Turbo solutions.
DAPO-Math-17k English Human + Model Construct Math Dataset of 17K diverse math problems for mathematical reasoning.
Big-Math-RL-Verified English Human + Model Construct Math Dataset of 251K math problems with verifiable answers, curated for reinforcement learning.
BELLE-math-zh Chinese Human + Model Construct Math Dataset of Chinese elementary school math problems with step-by-step solutions.
MathInstruct-Chinese Chinese Model Construct Math Chinese version of MathInstruct; contains instruction-style math problems in Chinese.
Legal-QA-v1 English Human Construct Law Dataset of 3.7K legal question-answer pairs sourced from legal forums.
Pile of Law English Human Construct Law Dataset for legal-domain tasks.
CUAD English Human Construct Law Expert-annotated dataset of 26K legal contract question-answer pairs across 41 clause categories.
LEDGAR English Human Construct Law Dataset of 1.45M contract clauses labeled by legal experts.
DISC-Law-SFT Chinese Human + Model Construct Law Comprehensive instruction tuning dataset covering diverse legal tasks.
Law-GPT-zh Chinese Human Construct Law Dataset of legal sentence pairs for training sentence embedding models in legal domain.
Lawyer LLaMA Data Chinese Human + Model Construct Law Dataset for Chinese legal tasks including legal consultation and bar exam question answering.
@@ 5.1 Instruction Fine-tuning Datasets
Table 9 presents a comprehensive collection of instruction fine-tuning datasets spanning several key domains, including general language understanding, finance, medicine, code, math, and law. For each domain, we select representative datasets from Hugging Face that are widely adopted by the community and closely aligned with real-world FedLLM applications. Building on this overview, we next provide a detailed introduction to the datasets within each domain.
@@ 5.1.1 General Instruction Fine-tuning Datasets
For English instruction fine-tuning, representative datasets include Alpaca (Taori et al., 2023), which is generated by Text-Davinci-003 with Alpaca-style instruction prompts, and Alpaca-GPT4 (Peng et al., 2023), which builds on this with GPT-4-generated, multi-turn dialogues to improve linguistic nuance and contextual coherence. Other commonly used datasets include Self-Instruct (Wang et al., 2022b), UltraChat 200k (Ding et al., 2023a), OpenOrca (Lian et al., 2023), ShareGPT-90K [^1] https://huggingface.co/datasets/liyucheng/ShareGPT90K, WizardLM Evol-Instruct V2 196k (Xu et al., 2023a), Databricks Dolly 15K (Conover et al., 2023), Baize (Xu et al., 2023b), OpenChat (Wang et al., 2023a), and Flan-v2222https://huggingface.co/datasets/SirNeural/flan__\__v2. A detailed comparison of these datasets is provided in Table 9.
@@
For Chinese instruction tuning, representative datasets include BELLE-train-0.5M-CN (Ji et al., 2023a), BELLE-train-1M-CN (Ji et al., 2023a), and BELLE-train-2M-CN (Ji et al., 2023a), which provide progressively scaled corpora designed to improve model alignment with Chinese user instructions. Additional resources such as Firefly-train-1.1M (Yang, 2023), Wizard-LM-Chinese-Instruct-Evol (Xu et al., 2023a), and HC3-Chinese (Guo et al., 2023a) support a wide range of instruction tasks, including reasoning, question answering, and domain-specific knowledge modeling.
@@
Bilingual instruction datasets are also included to support multilingual and cross-lingual instruction tuning in federated contexts. HC3 (Guo et al., 2023a) offers both human- and model-generated responses in English and Chinese for evaluating factual consistency and detection capabilities. ShareGPT-Chinese-English-90k (shareAI, 2023) provides high-quality, bilingual conversations, making it particularly suitable for instruction tuning in multilingual FedLLM scenarios.

### v3

@@ 4.1.2 LoRA in Federated Fine-Tuning
FedIT (Zhang et al., 2024b) directly integrates LoRA into the classic FedAvg (McMahan et al., 2017) for instruction tuning. FedSA-LoRA (Guo et al., 2024b) identifies that the $\mathbf{A}$ matrices primarily encode general knowledge, while the $\mathbf{B}$ matrices capture client-specific features; thus, it only uploads $\mathbf{A}$ to the server, significantly reducing communication overhead. FederatedScope-LLM (Kuang et al., 2023) establishes a comprehensive end-to-end pipeline for federated LLM fine-tuning and proposes offsite-tuning strategies to mitigate both communication and computational costs. FeDeRA (Yan et al., 2024) addresses data heterogeneity by initializing LoRA matrices via singular value decomposition on the pre-trained weights. LoRA-FAIR (Bian et al., 2024) introduces a correction mechanism on the server to handle aggregation bias and initialization drift across clients. FLASC (Kuo et al., 2024) incorporates sparsity into LoRA to further reduce communication overhead. SA-FedLoRA (Yang et al., 2024c) mitigates client drift through parameter regularization and dynamically allocates communication budgets.
SLoRA (Babakniya et al., 2023) proposes a novel data-driven initialization scheme to better handle statistical heterogeneity. RoLoRA (Chen et al., 2024d) adopts an alternating minimization approach to improve robustness under non-IID conditions. FedPipe (Fang et al., 2024b) automatically selects critical parameters for fine-tuning and applies quantization to reduce memory usage. LP-FL (Jiang et al., 2023) applies LoRA directly to enable efficient on-device fine-tuning. Fed-piLot (Zhang et al., 2024l) reduces memory consumption through LoRA assignment strategies and introduces a novel spatial-temporal aggregation (STAgg) rule to address heterogeneity. FedRA (Su et al., 2023) adaptively determines parameter update scopes based on client resource constraints, effectively reducing computational, communication, and memory costs. FedPruner (Wu et al., 2025c) introduces a macro-micro synergetic pruning framework to mitigate memory constraints on participating devices. To further facilitate large-scale deployment, DevFT (Wu et al., 2025b) proposes a developmental federated tuning paradigm to minimize training resource overheads.
@@
FedPruner (Wu et al., 2025c)
DevFT (Wu et al., 2025b)
@@
SmartFed (Wu et al., 2025a)
$\bullet$ Heterogeneous LoRA allows clients to adopt different rank values $r$ based on their data characteristics or resource constraints. This heterogeneity can manifest either across clients (inter-model) or within different layers of the same model (intra-model). By enabling each client to select a rank that best fits its capabilities and local data, this approach introduces greater flexibility and resource-awareness into the federated fine-tuning process. Table 3 summarizes representative methods and the specific challenges they address.
HETLoRA (Cho et al., 2024) assigns heterogeneous ranks across devices and incorporates rank self-pruning along with sparsity-weighted aggregation to tackle data heterogeneity. FLoRA (Wang et al., 2024f) proposes a stacking-based aggregation scheme and allows devices to select ranks according to their resource budgets. FlexLoRA (Bai et al., 2024a) enables dynamic adjustment of local LoRA ranks to leverage the heterogeneous device resources, while employing singular value decomposition for weight redistribution. LoRA-A2 (Koo et al., 2024) introduces alternating freezing and adaptive rank selection mechanisms to fully utilize heterogeneous device resources while addressing statistical heterogeneity. Byun & Lee (2024) propose a replication-based padding technique to enable aggregation across clients with varying LoRA ranks. FEDHM (Yao et al., 2021) addresses resource constraints by distributing low-rank models with heterogeneous capacities to clients. RBLA (Tavallaie & Nazemi1, ) improves aggregation robustness by simultaneously maintaining and aligning both low-rank and high-rank feature components. SmartFed (Wu et al., 2025a) achieves efficient task adaptation by leveraging rank-wise reconfiguration of existing LoRA modules, thereby avoiding the prohibitive computational overhead of training from scratch.
@@
FDLoRA (Qi et al., 2024a) introduces dual LoRA modules on each client to separately capture global and personalized knowledge. pFedLoRA (Yi et al., 2023) designs a homogeneous small adapter to facilitate federated clients’ heterogeneous local model training, with a proposed iterative training process for global-local knowledge exchange. FedLoRA (Wu et al., 2024f) maintains shared general knowledge in a global full-rank matrix while encoding client-specific knowledge in a personalized low-rank module. FedDPA (Yang et al., 2024b) utilizes a global adapter and a local adapter to jointly address test-time distribution shifts and client-specific personalization. PerFIT (Zhang et al., 2024e) allows each client to search for a personalized architecture by expanding the trainable parameter space of the global model to address data heterogeneity. FEDMEM (Du et al., 2024) equips the global model with a KNN classifier that captures client-specific distributional shifts, achieving personalization and overcoming data heterogeneity. FedAMoLE (Zhang et al., 2024i) features a mixture of LoRA experts module for aggregating heterogeneous models and a reverse selection-based expert assignment strategy that optimizes model architectures based on data distributions.
@@ 5.1 Instruction Fine-Tuning Datasets
Table 9 and Table 10 present a comprehensive collection of instruction fine-tuning datasets spanning several key domains, including general language understanding, finance, medicine, code, math, and law. For each domain, we select representative datasets from Hugging Face that are widely adopted by the community and closely aligned with real-world FedLLM applications. Building on this overview, we next provide a detailed introduction to the datasets within each domain.
@@ 5.1.1 General Instruction Fine-Tuning Datasets
For English instruction fine-tuning, representative datasets include Alpaca (Taori et al., 2023), which is generated by Text-Davinci-003 with Alpaca-style instruction prompts, and Alpaca-GPT4 (Peng et al., 2023), which builds on this with GPT-4-generated, multi-turn dialogues to improve linguistic nuance and contextual coherence. Other commonly used datasets include Self-Instruct (Wang et al., 2022b), UltraChat 200k (Ding et al., 2023a), OpenOrca (Lian et al., 2023), ShareGPT-90K [^1] https://huggingface.co/datasets/liyucheng/ShareGPT90K, WizardLM Evol-Instruct V2 196k (Xu et al., 2023a), Databricks Dolly 15K (Conover et al., 2023), Baize (Xu et al., 2023b), OpenChat (Wang et al., 2023a), and Flan-v2222https://huggingface.co/datasets/SirNeural/flan_\_v2.
@@
For Chinese instruction tuning, the landscape includes representative datasets such as BELLE-CN (Ji et al., 2023a), which focuses on improving model alignment with Chinese user instructions. This is complemented by additional resources like Firefly-train-1.1M (Yang, 2023), Wizard-LM-Chinese-Instruct-Evol (Xu et al., 2023a), and HC3-Chinese (Guo et al., 2023a). These resources facilitate a diverse range of instruction tasks, including reasoning, question answering, and domain-specific knowledge modeling.
@@
Bilingual instruction datasets are also included to support multilingual and cross-lingual instruction tuning in federated contexts. HC3 (Guo et al., 2023a) offers both human- and model-generated responses in English and Chinese for evaluating factual consistency and detection capabilities. ShareGPT-Chinese-English-90k (shareAI, 2023) provides high-quality, bilingual conversations, making it particularly suitable for instruction tuning in multilingual FedLLM scenarios. A detailed comparison is provided in Table 9.
Table 9: A summary of representative instruction fine-tuning datasets in general, financial, and medical domains. The construction methods are categorized into three types: human construct, which refers to datasets written or annotated by humans; model construct, which refers to data generated by prompting LLMs; and synthetic, which refers to data produced through rule-based or programmatic generation.
Dataset Language Construction Domain Description
Alpaca English Model General Generated by Text-Davinci-003 (Alpaca-style)
Alpaca-GPT4 English Model General Generated by GPT-4 (multi-turn Alpaca prompts)
Self-Instruct English Human + Model General Generated from seed instructions via GPT-3
UltraChat 200k English Model General Multi-turn dialogues filtered from UltraChat
OpenOrca English Model General  4.2M GPT-3.5/4 augmented FLAN examples
ShareGPT90K English Model General 90K multi-turn dialogues from ShareGPT
WizardLM Evol-Instruct V2 English Model General 196K samples generated via Evol-Instruct
Databricks Dolly 15K English Human General 15K human-generated prompt-response pairs
Baize English Model General Dialogues from ChatGPT (prompted by user queries)
OpenChat English Model General Dialogues from open LLMs for multi-turn alignment
Flan-v2 English Model General Combination of Flan, P3, SNI, CoT, and Dialog tasks
BELLE-CN Chinese Human + Model General 2M Chinese instruction-following samples
Firefly-train-1.1M Chinese Human General 1.65M Chinese samples (23 tasks, human templates)
Wizard-LM-Chinese-instruct-evol Chinese Human + Model General 70K Chinese samples (translated from WizardLM)
HC3-Chinese Chinese Human + Model General Chinese Human-ChatGPT QA pairs (multi-domain)
HC3 English / Chinese Human + Model General Human-ChatGPT QA pairs (multi-domain)
ShareGPT-Chinese-English-90k English / Chinese Model General 90K bilingual human-machine QA pairs
FinGPT English Human + Model Financial Instruction-tuning data for financial tasks
Finance-Instruct-500k English Human + Model Financial Large-scale instruction-tuning data for financial tasks
Finance-Alpaca English Human + Model Financial Instruction-tuning data combining Alpaca and FiQA
Financial PhraseBank English Human Financial Manually annotated financial news for sentiment
Yahoo-Finance-Data English Human Financial Yahoo finance data
Financial-QA-10K English Model Financial Contextual QA for financial answering and retrieval
Financial-Classification English Human Financial Financial PhraseBank and Kaggle financial texts
Twitter-Financial-News-Topic English Human Financial Annotated tweets for financial topic classification
Financial-News-Articles English Human Financial Financial articles for classification and sentiment analysis
FiQA English Human Financial QA from financial texts and forums
Earnings-Call English Human Financial QA pairs from CEO/CFO earnings calls
Doc2EDAG Chinese Human Financial Financial reports annotated for event graph extraction
Synthetic-PII-Finance-Multilingual Multilingual Synthetic Financial Synthetic financial docs with labeled PII
ChatDoctor-200K English Human + Model Medical Medical QA and dialogue instructions
ChatDoctor-HealthCareMagic-100k English Human Medical Real-world doctor-patient conversations
Medical Meadow CORD-19 English Human Medical Literature summarization instructions (CORD-19)
Medical Meadow MedQA English Human Medical Medical QA from professional board exams
HealthCareMagic-100k-en English Human Medical Medical and real doctor-patient consultations
ChatMed-Consult-Dataset Chinese Human + Model Medical Medical consultation QA and dialogue instructions
CMtMedQA Chinese Human Medical Multi-turn real medical QA
DISC-Med-SFT Chinese Human + Model Medical Real dialogues and knowledge graph QA pairs
HuatuoGPT-SFT Chinese Human + Model Medical ChatGPT-generated and real doctor-patient dialogues
Huatuo26M-Lite Chinese Human + Model Medical Refined subset of Huatuo-26M (ChatGPT-rewritten)
ShenNong-TCM Chinese Human + Model Medical QA pairs from TCM knowledge graph
MedDialog English / Chinese Human Medical Large-scale doctor-patient dialogues